OpenStudy (anonymous):

Find the sum of the first 34 terms of the arithmetic sequence whose first term is -8 and whose 34th term is -173.

OpenStudy (anonymous):

Sn = n/2(2a + (n-1)d) so S34 = 17(-16 + 33*d) d remains to be found nth term is a + (n-1)d, so -173 = -8 + 33d, so 33d = -165, so d = -5 so S34 = 17(-16 - 165) = 17*-181